# A Theoretical Framework for Data Efficient Multi-Source Transfer Learning Based on Cramér-Rao Bound

### v2

@@ Appendix B Proofs
In this section, for the sake of clarity, we will write $\hat{\theta}$ in its parameterized form $\hat{\theta}(X^{N_{0}})$ when necessary, and these two forms are mathematically equivalent. By taking Taylor expansion of $D\left(P_{X;\hat{\theta}(X^{N_{0}})}\middle\|P_{X;\theta_{0}}\right)$ at $\theta_{0}$, we can get
$D\left(P_{X;\hat{\theta}(X^{N_{0}})}\middle\|P_{X;\theta_{0}}\right)=\frac{1}{2}\sum_{x\in X}\frac{\left(P_{X;\hat{\theta}(X^{N_{0}})}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})$ (17)
@@
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}(X^{N_{0}})}\middle\|P_{X;\theta_{0}}\right)\right]$
$=\sum_{X^{N_{0}}}P_{X^{n};\theta_{0}}(X^{N_{0}})D\left(P_{X;\hat{\theta}(X^{N_{0}})}\middle\|P_{X;\theta_{0}}\right)$
$=\sum_{X^{N_{0}}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(P_{X;\hat{\theta}(X^{N_{0}})}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})\right)$
$=\sum_{X^{N_{0}}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})\right)$
=$=\sum_{\left\{X^{N_{0}}:\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}<\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})\right)$
+$+\sum_{\left\{X^{N_{0}}:\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}\geq\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})\right)$
To facilitate the subsequent proof, we introduce the concept of ”Dot Equal”.
Definition B.2.
(Dot Equal ($\dot{=}$)) Specifically, given two functions $f(n)$ and $g(n)$, the notation $f(n)\dot{=}g(n)$ is defined as
@@
We denote $\hat{P}_{X^{N_{0}}}$ as the empirical distribution of $X^{N_{0}}$. Applying Sanov’s Theorem to (B.1), we can know that
@@
Then, we aim to establish a connection between the K-L divergence (20) and $\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}$. From (17), we can know that
$D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)=\frac{1}{2}\sum_{x\in X}\frac{\left(\hat{P}_{X^{N_{0}}}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})$ (21)
From the features of MLE, we can know that
$\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial\log P_{X;{\hat{\theta}(X^{N_{0}})}}(x)}{\partial\hat{\theta}}\right]=0=\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}\right]+\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial^{2}\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}^{2}}\right]{\left(\hat{\theta}(X^{N_{0}})-\theta_{0}\right)},$ (22)
@@
$\left(\hat{\theta}(X^{N_{0}})-\theta_{0}\right)=\frac{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}\right]}{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial^{2}\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}^{2}}\right]}=\frac{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\frac{\partial P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}}{P_{X;{\theta_{0}}}(x)}\right]}{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial^{2}\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}^{2}}\right]}=\frac{\sum\limits_{x\in\mathcal{X}}{\left(\hat{P}_{X^{N_{0}}}(x)-P_{X;\theta_{0}}(x)\right)}\frac{\frac{\partial P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}}{P_{X;{\theta_{0}}}(x)}}{J(\theta_{0})}$ (23)
@@
Combining with (21), (23), and (25), the inequality (24) can be transformed to
$D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)\geq\frac{J(\theta_{0})\left(\hat{\theta}(X^{N_{0}})-\theta_{0}\right)^{2}}{2}=\frac{J(\theta_{0})\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}}{2}$ (26)
Combining (20) and (26), we can know that
$P_{X^{n};\theta_{0}}(X^{N_{0}})\doteq e^{-N_{0}D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)}\leq e^{\frac{-N_{0}J(\theta_{0})\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}}{2}}$ (27)
For the first term in (B.1), the $\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|$ is small enough for us to omit the term $o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})$. As for the second term in in (B.1), even though the magnitude of $\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|$ is no longer negligible, the probability of such sequences is $O(e^{\frac{-N_{0}J(\theta_{0})\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}}{2}})$ by (27), which is exponentially decaying with $N_{0}$ such that the second term is $o{(\frac{1}{N_{0}})}$. By transfering (B.1), we can get
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}(X^{N_{0}})}\middle\|P_{X;\theta_{0}}\right)\right]$
=$=\sum_{\left\{X^{N_{0}}:\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}<\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right)+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})$
[…]
$=\frac{1}{2}\mathbb{E}\left[\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right]+o{(\frac{1}{N_{0}})}.$ (28)
We then transform (B.1) with (25)
@@
Combining (B.1), (B.1), we can get
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}}\middle\|P_{X;\theta_{0}}\right)\right]=\frac{1}{2}\mathbb{E}\left[\left(\hat{\theta}-\theta_{0}\right)^{2}\right]{J}(\theta_{0})+o{(\frac{1}{N_{0}})}.$ (30)
From (B.1), we can establish the relationship between the proposed K-L measure (6) and the mean squared error. Then, by using the (3), the K-L measure is
@@
B.2 Proof of Theorem 4.3
In this section we will be using samples from both distributions for our maximum likelihood estimation, so our optimization problem becomes
@@
Lemma B.3.
@@
Proof.
Since $\hat{\theta}$ is a maximizer of $L_{n}{(\theta)}$, $L_{n}^{{}^{\prime}}{(\hat{\theta})}=0=L_{n}^{{}^{\prime}}{(E_{\hat{\theta}})}+L_{n}^{{}^{\prime\prime}}{(E_{\hat{\theta}})}(\hat{\theta}-E_{\hat{\theta}})$. Therefore,
$\sqrt{N_{0}+n_{1}}(\hat{\theta}-E_{\hat{\theta}})=\frac{\sqrt{N_{0}+n_{1}}L_{n}^{{}^{\prime}}{(E_{\hat{\theta}})}}{L_{n}^{{}^{\prime\prime}}{(E_{\hat{\theta}})}}.$ (39)
Since $E_{\hat{\theta}}$ maximizes $L{(\theta)}$, $L^{{}^{\prime}}{(E_{\hat{\theta}})}=\frac{N_{0}}{N_{0}+n_{1}}\mathbb{E}_{\theta_{0}}\left[\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial x}\right]+\frac{n_{1}}{N_{0}+n_{1}}\mathbb{E}_{\theta_{1}}\left[\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x_{i}^{{}^{\prime}})}}{\partial x}\right]=0$. Therefore,
$\sqrt{N_{0}+n_{1}}L_{n}^{{}^{\prime}}(E_{\hat{\theta}})$
@@
Applying the Central Limit Theorem to (B.2), we can get
@@
Under the assumption that $\theta_{0},\theta_{1},E_{\hat{\theta}}$ are sufficiently close to each other, we can easily deduce that the difference among ${J}(\theta_{0}),{J}(\theta_{1})$ and ${J}(E_{\hat{\theta}})$ is $O(\frac{1}{\sqrt{N_{0}+n_{1}}})$. We can easily get
@@
Therefore, the bound of $\mathbb{E}\left[\left(\hat{\theta}-E_{\hat{\theta}}\right)^{2}\right]$ is
@@
Combining (6) and (B.1), we know that
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}}\middle\|P_{X;\theta_{0}}\right)\right]$
@@
Combining (B.2), (37) and (47), we know that the K-L measure is
$\frac{1}{2}\frac{1}{N_{0}+n_{1}}+\frac{1}{2}{J}(\theta_{0})\frac{n_{1}^{2}}{(N_{0}+n_{1})^{2}}\left(\theta_{1}-\theta_{0}\right)^{2}$ (49)
B.3 Proof of Proposition 4.4
Similar to (B.1), we can get
$\mathbb{E}\left[D\left(P_{X;\hat{{\underline{\theta}}}}\middle\|P_{X;{\underline{\theta}}_{0}}\right)\right]$

### v3

@@ C.1 Proof of Lemma 3
In this section, for the sake of clarity, we will write $\hat{\theta}$ in its parameterized form $\hat{\theta}(X^{N_{0}})$ when necessary, and these two forms are mathematically equivalent. By taking Taylor expansion of $D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}(X^{N_{0}})}\right)$ at $\theta_{0}$, we can get
$D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}(X^{N_{0}})}\right)$
$=\sum_{x\in X}P_{X;\theta_{0}}(x)\log\frac{P_{X;\theta_{0}}(x)}{P_{X;\hat{\theta}(X^{N_{0}})}(x)}$
$=-\sum_{x\in X}P_{X;\theta_{0}}(x)\log\frac{P_{X;\hat{\theta}(X^{N_{0}})}(x)}{P_{X;\theta_{0}}(x)}$
$=-\sum_{x\in X}P_{X;\theta_{0}}(x)\log\left(1+\frac{P_{X;\hat{\theta}(X^{N_{0}})}(x)-P_{X;\theta_{0}}(x)}{{P_{X;\theta_{0}}(x)}}\right)$
$=-\sum_{x\in X}P_{X;\theta_{0}}(x)\left(\left(\frac{P_{X;\hat{\theta}(X^{N_{0}})}(x)-P_{X;\theta_{0}}(x)}{{P_{X;\theta_{0}}(x)}}\right)-\frac{1}{2}\left(\frac{P_{X;\hat{\theta}(X^{N_{0}})}(x)-P_{X;\theta_{0}}(x)}{{P_{X;\theta_{0}}(x)}}\right)^{2}\right)+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$
$=\frac{1}{2}\sum_{x\in X}\frac{\left(P_{X;\hat{\theta}(X^{N_{0}})}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$ (17)
@@
$\mathbb{E}\left[D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}(X^{N_{0}})}\right)\right]$
$=\sum_{X^{N_{0}}}P_{X^{n};\theta_{0}}(X^{N_{0}})D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}(X^{N_{0}})}\right)$
$=\sum_{X^{N_{0}}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(P_{X;\hat{\theta}(X^{N_{0}})}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})\right)$
$=\sum_{X^{N_{0}}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})\right)$
$=\sum_{\left\{X^{N_{0}}:|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}<\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})\right)$
$+\sum_{\left\{X^{N_{0}}:|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}\geq\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})\right)$
To facilitate the subsequent proof, we introduce the concept of "Dot Equal".
Definition 9.
[…]
@@
We denote $\hat{P}_{X^{N_{0}}}$ as the empirical distribution of $X^{N_{0}}$. Applying Sanov’s Theorem to (C.1), we can know that
@@
Then, we aim to establish a connection between (20) and $|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}$. From (C.1), we can know that the $D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)$ in (20) can be transformed to
$D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)=\frac{1}{2}\sum_{x\in X}\frac{\left(\hat{P}_{X^{N_{0}}}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{\hat{P}_{X^{N_{0}}}(x)}+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$
$=\frac{1}{2}\sum_{x\in X}\frac{\left(\hat{P}_{X^{N_{0}}}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$ (21)
From the characteristics of MLE, we can know that
$\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial\log P_{X;{\hat{\theta}(X^{N_{0}})}}(x)}{\partial\hat{\theta}}\right$]
$=0$
$=\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}\right]+\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial^{2}\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}^{2}}\right]{\left(\hat{\theta}(X^{N_{0}})-\theta_{0}\right)}+O(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}),$ (22)
@@
$\left(\hat{\theta}(X^{N_{0}})-\theta_{0}\right)+O(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$
$=-\frac{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}\right]}{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial^{2}\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}^{2}}\right]}$
$=-\frac{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\frac{\partial P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}}{P_{X;{\theta_{0}}}(x)}\right]}{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial^{2}\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}^{2}}\right]}$
$=\frac{\sum\limits_{x\in\mathcal{X}}{\left(\hat{P}_{X^{N_{0}}}(x)-P_{X;\theta_{0}}(x)\right)}\frac{\frac{\partial P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}}{P_{X;{\theta_{0}}}(x)}}{J(\theta_{0})}$ (23)
@@
Combining with (C.1), (C.1), and (25), the inequality (24) can be transformed to
$D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)$
$=\frac{1}{2}\sum_{x\in X}\frac{\left(\hat{P}_{X^{N_{0}}}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$
$\geq\frac{1}{2}J(\theta_{0})\left(\hat{\theta}(X^{N_{0}})-\theta_{0}+O(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})\right)^{2}+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$
$=\frac{1}{2}J(\theta_{0})|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$ (26)
Combining (20) and (C.1), we can know that
$P_{X^{n};\theta_{0}}(X^{N_{0}})\doteq e^{-N_{0}D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)}\leq e^{\frac{-N_{0}J(\theta_{0})|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}}{2}}$ (27)
For the first term in (C.1), the $\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|$ is small enough for us to omit the term $o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$. As for the second term in in (C.1), even though the magnitude of $\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|$ is no longer negligible, the probability of such sequences is $O(e^{\frac{-N_{0}J(\theta_{0})|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}}{2}})$ by (27), which is exponentially decaying with $N_{0}$ such that the second term is $o{(\frac{1}{N_{0}})}$. By transfering (C.1), we can get
$\mathbb{E}\left[D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}(X^{N_{0}})}\right)\right]$
$=\sum_{\left\{X^{N_{0}}:|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}<\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right)+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$
$+\sum_{\left\{X^{N_{0}}:|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}\geq\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right)+o{(\frac{1}{N_{0}})}$
$=\frac{1}{2}\mathbb{E}\left[\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right]+o{(\frac{1}{N_{0}})}.$ (28)
We then transform (C.1) with (25)
@@
Combining (C.1), (C.1), we can get
$\mathbb{E}\left[D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}}\right)\right]=\frac{1}{2}\mathbb{E}\left[\left(\hat{\theta}-\theta_{0}\right)^{2}\right]{J}(\theta_{0})+o{(\frac{1}{N_{0}})}.$ (30)
From (8), we can establish the relationship between the proposed proposed measure (6) and the mean squared error. Then, by using the (2), the K-L measure is
@@
C.2 Proof of Theorem 4
In this theorem, we are using samples from both target task and source task for our maximum likelihood estimation, so our optimization problem becomes
@@ C.2 Proof of Theorem 4
where $L(\theta)$ denotes the expectation of $L_{n}(\theta)$
$L(\theta)\triangleq\frac{N_{0}}{N_{0}+n_{1}}\mathbb{E}_{P_{X;\theta_{0}}}\left[\log P_{X;\theta}(x)\right]+\frac{n_{1}}{N_{0}+n_{1}}\mathbb{E}_{P_{X;\theta_{1}}}\left[\log P_{X;\theta}(x)\right].$ (35)
@@
Proof.
From (36), we know that
$E_{\hat{\theta}}=\operatorname*{arg\,min}_{\theta}\frac{N_{0}D\left(P_{X;\theta_{0}}\middle\|P_{X;\theta}\right)}{N_{0}+n_{1}}+\frac{n_{1}D\left(P_{X;\theta_{1}}\middle\|P_{X;\theta}\right)}{N_{0}+n_{1}}$
$=\operatorname*{arg\,min}_{\theta}\frac{N_{0}}{N_{0}+n_{1}}\sum_{x\in\mathcal{X}}P_{X;\theta_{0}}(x)\log\frac{P_{X;\theta_{0}}(x)}{P_{X;\theta}(x)}+\frac{n_{1}}{N_{0}+n_{1}}\sum_{x\in\mathcal{X}}P_{X;\theta_{1}}(x)\log\frac{P_{X;\theta_{1}}(x)}{P_{X;\theta}(x)}$ (39)
To minimize the weighted K-L divergence. We treat $P_{X;\theta}(x),\forall x\in\mathcal{X}$ as the variable with the constraint $\sum\limits_{x}P_{X;\theta}(x)=1$, then we form the Lagrangian:
$\text{Lagrangian}(P,\lambda)=-\sum_{x}\left(\frac{N_{0}}{N_{0}+n_{1}}P_{X;\theta_{0}}(x)+\frac{n_{1}}{N_{0}+n_{1}}P_{X;\theta_{1}}(x)\right)\log P_{X;\theta}(x)+\lambda\left(\sum_{x}P_{X;\theta}(x)-1\right)$ (40)
Taking the derivative with respect to $P_{X;\theta}(x)$ and setting it to zero gives:
$\frac{\partial\text{Lagrangian}(P,\lambda)}{\partial P_{X;\theta}(x)}=-\frac{\frac{N_{0}}{N_{0}+n_{1}}P_{X;\theta_{0}}(x)+\frac{n_{1}}{N_{0}+n_{1}}P_{X;\theta_{1}}(x)}{P_{X;\theta}(x)}+\lambda=0.$ (41)
So
$P_{X;\theta}(x)=\frac{\frac{N_{0}}{N_{0}+n_{1}}P_{X;\theta_{0}}(x)+\frac{n_{1}}{N_{0}+n_{1}}P_{X;\theta_{1}}(x)}{\lambda}.$ (42)
Normalizing $P_{X;\theta}(x)$ gives $\lambda=1$, hence the optimal solution is:
$P_{X;E_{\hat{\theta}}}(x)=\frac{N_{0}}{N_{0}+n_{1}}P_{X;\theta_{0}}(x)+\frac{n_{1}}{N_{0}+n_{1}}P_{X;\theta_{1}}(x),\forall x\in\mathcal{X},$ (43)
which corresponds to (37).
Then we begin to prove (38). By doing a Taylor Expansion of (43) around $\theta_{0}$, we can get
$P_{X;\theta_{0}}(x)+\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta}(E_{\hat{\theta}}-\theta_{0})+O(|E_{\hat{\theta}}-\theta_{0}|^{2})$
$=\frac{N_{0}}{N_{0}+n_{1}}P_{X;\theta_{0}}(x)+\frac{n_{1}}{N_{0}+n_{1}}\left(P_{X;\theta_{0}}(x)+\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta}(\theta_{1}-\theta_{0})+O(|\theta_{0}-\theta_{1}|^{2})\right)$ (44)
From (C.2) we can get
$E_{\hat{\theta}}=\frac{N_{0}\theta_{0}+n_{1}\theta_{1}}{N_{0}+n_{1}}+O(|\theta_{0}-\theta_{1}|^{2}),$ (45)
So we can get
$E_{\hat{\theta}}=\frac{N_{0}\theta_{0}+n_{1}\theta_{1}}{N_{0}+n_{1}}+O\left(\frac{1}{N_{0}+n_{1}}\right).$ (46)
Lemma 11.
We assume that the following regularity conditions hold:
1. The log-likelihood function is twice continuously differentiable in the neighborhood of $\theta_{0}$.
2. The Fisher information $J(\theta_{0})$ is positive and finite.
Then, the estimator $\hat{\theta}$ is asymptotically normal,i.e.,
$\sqrt{N_{0}+n_{1}}(\hat{\theta}-E_{\hat{\theta}})\xrightarrow{d}\mathcal{N}\left(0,\frac{1}{J(\theta_{0})}\right).$ (47)
Proof.
Since $\hat{\theta}$ is a maximizer of $L_{n}{(\theta)}$, $L_{n}^{{}^{\prime}}{(\hat{\theta})}=0=L_{n}^{{}^{\prime}}{(E_{\hat{\theta}})}+L_{n}^{{}^{\prime\prime}}{(E_{\hat{\theta}})}(\hat{\theta}-E_{\hat{\theta}})+O\left(\frac{1}{N_{0}}\right)$. Therefore,
$\sqrt{N_{0}+n_{1}}(\hat{\theta}-E_{\hat{\theta}})=-\frac{\sqrt{N_{0}+n_{1}}L_{n}^{{}^{\prime}}{(E_{\hat{\theta}})}}{L_{n}^{{}^{\prime\prime}}{(E_{\hat{\theta}})}}+O\left(\frac{1}{\sqrt{N_{0}}}\right).$ (48)
Since $E_{\hat{\theta}}$ maximizes $L{(\theta)}$, $L^{{}^{\prime}}{(E_{\hat{\theta}})}=\frac{N_{0}}{N_{0}+n_{1}}\mathbb{E}_{\theta_{0}}\left[\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial x}\right]+\frac{n_{1}}{N_{0}+n_{1}}\mathbb{E}_{\theta_{1}}\left[\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x_{i})}}{\partial x}\right]=0$. Therefore,
@@
$=\sqrt{\frac{N_{0}}{N_{0}+n_{1}}}\left(\sqrt{\frac{1}{N_{0}}}\sum_{x\in X^{N_{0}}}\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)+\sqrt{\frac{n_{1}}{N_{0}+n_{1}}}\left(\sqrt{\frac{1}{n_{1}}}\sum_{x\in X^{n_{1}}}\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)$
$=\sqrt{\frac{N_{0}}{N_{0}+n_{1}}}\left(\sqrt{\frac{1}{N_{0}}}\sum_{x\in X^{N_{0}}}\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)+\sqrt{\frac{n_{1}}{N_{0}+n_{1}}}\left(\sqrt{\frac{1}{n_{1}}}\sum_{x\in X^{n_{1}}}\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)$
$-\frac{N_{0}}{\sqrt{{N_{0}+n_{1}}}}\mathbb{E}_{\theta_{0}}\left[\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial x}\right]-\frac{n_{1}}{\sqrt{{N_{0}+n_{1}}}}\mathbb{E}_{\theta_{1}}\left[\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial x}\right]$
@@
Applying the Central Limit Theorem to (C.2), we can get
@@
$\sqrt{N_{0}+n_{1}}(\hat{\theta}-E_{\hat{\theta}})\xrightarrow{d}\mathcal{N}\left(0,\frac{1}{J(\theta_{0})}\right)$ (55)
Therefore, the limit of $\mathbb{E}\left[\left(\hat{\theta}-E_{\hat{\theta}}\right)^{2}\right]$ is
@@
Combining (6) and (C.1), we know that
[…]
@@
Combining (C.2), (38) and (56), we know that the proposed measure is
$\frac{1}{2}\frac{1}{N_{0}+n_{1}}+\frac{1}{2}{J}(\theta_{0})\frac{n_{1}^{2}}{(N_{0}+n_{1})^{2}}\left(\theta_{1}-\theta_{0}\right)^{2}+o(\frac{1}{N_{0}})$ (58)
C.3 Proof of Proposition 5
Similar to (8), we can get
[…]